What is the equation for a circle with a diameter having the following endpoints: N(0,1) and Q(18,1)?
Answer by user_dude2008(1862)

Center:

(0+18)/2=9, (1+1)/2=1

Center is (9,1)


Radius:

r=sqrt((0-18)^2+(1-1)^2)/2 = 9

radius is 9


Ans: Equation is